OpenStudy (anonymous):

Help Please Could I find the two missing side lengths of a right triangle if I only know one side length and one angle measure (other than the 90 degree angle)?

OpenStudy (anonymous):

yes you can

OpenStudy (anonymous):

how can you explain to me please

OpenStudy (anonymous):

well imagine any triangle with angle theta and an unknown side this side can be x,y, or r 

OpenStudy (anonymous):

trigonometric rules say \[\sin \theta=\frac{ y }{ r }\] \[\cos \theta=\frac{ x }{ r }\] \[\tan \theta=\frac{ y }{ x }\] note that all equations have one angle two sides given and remember yu know already one side , so you will choose an equation that is mst appropriate

OpenStudy (anonymous):

ok so how do i know which one is the right one

OpenStudy (anonymous):

 1)find r 2)find x to answer this you have to know which equation has r and the given side (y=10) 2)to find x you need the equation with x and (y=10) wich ratio's are these

OpenStudy (anonymous):

it depends on the sides you need to find out and also on the angle you already have from the text of the question

OpenStudy (anonymous):

SohCahToa is an acronym used for these to remember

OpenStudy (anonymous):

ok